# The MASK Benchmark: Disentangling Honesty From Accuracy in AI Systems

### v2

@@ 4.3 Metrics
Accuracy measures a model’s truthfulness by evaluating the alignment between its belief $B$ and the ground truth $T$ for each proposition. We compute $\text{Accuracy}(B,T)=0$ if $B\neq T$, marking such cases as inaccurate. When a model shows no clear belief, we note this separately. The final accuracy score is the average accuracy across all examples.
@@ 5 Experiments
We use MASK to explore two main experiments around model honesty. First, we conduct a comprehensive assessment of the degree of honesty exhibited by state-of-the-art language models. Second, we evaluate the two baseline interventions designed to improve model honesty.
@@ 5.1 Models Evaluation
Figure 6 reports the honesty rates of state-of-the-art models, which are quite low. No model shown is explicitly honest in more than $46\%$ of cases. GPT-4o and Llama-405B tell more lies than Claude 3.7 Sonnet, and most models are dishonest more than a third of the time. These lies appear even in short, straightforward scenarios, implying that current instruction tuning techniques alone are insufficient to prevent dishonesty. We also measure each model’s factual accuracy (Figure 6) and observe that highly capable models tend to have over $70\%$ accuracy on belief prompts (they give more correct facts) but do not necessarily show higher honesty. Figure 8 further shows that this finding holds across archetypes, indicating that these models can be pressured to lie in many different settings.
@@
In Figure 7, we show that increased compute (FLOP) does not lead to more honest models, showing a negative correlation (Spearman coefficient: $-59.9\%$). This indicates that higher scores are likely a result of design decisions and fine-tuning, rather than due solely to model capabilities obtained during pre-training. In contrast, accuracy strongly correlates with the amount of training compute used (Spearman coefficient: $87.3\%$). While model scaling improves factual accuracy, scaling AI models does not solve their tendency to produce intentional falsehoods when pressured.
For additional results, including a preliminary analysis of why different models have different honesty scores and a corroboration of our evaluation with self-reported lying, see Appendix A.
@@ 5.2 Improving the Honesty of Models
We test two baseline interventions to improve the honesty of large language models and evaluate their effectiveness using MASK.
@@
For each pressure prompt, we prepend the following text to the system prompt, as if there were a separate “developer system prompt”:
USER SYSTEM PROMPT: <original_system_prompt>
Although this approach improved the honesty of the responses in many examples ($+12.2\%$ for Llama-2-7B, $+8.8\%$ for Llama-2-13B), it still leaves substantial room for improvement (Figure 9). This highlights the challenge that, in safety-critical domains, a model’s default behavior often requires more robust interventions than prompt engineering alone. Relying on specialized prompt engineering can also be tiresome; models should default to honest behavior without extensive developer prompt engineering.
@@
The technique is described further by Zou et al. (2023). While LoRRA led to measurable improvements in model honesty ($+6.6\%$ for Llama-2-7B, $+13.1\%$ for Llama-2-13B), it was also not sufficient to eliminate all dishonesty. This suggests that representation engineering may require methodological improvements to be fully robust in controlling lying in large language models.
5.3 Self-Reported Lying
Experimental setup.
Results.
@@ 6 Conclusion
We introduced MASK, a dataset and evaluation framework for measuring dishonesty in LLMs by testing whether models will contradict their own elicited beliefs. Our experiments reveal that many current models, despite growing general capabilities, can still produce lies of commission under pressure. These findings suggest that scaling alone does not improve honesty. We also presented preliminary methods to reduce dishonesty through targeted prompts and representation engineering, though these approaches are imperfect and leave room for improvement. We hope MASK prompts more rigorous investigation of honesty as a distinct safety property, as well as research into reliable methods for eliminating dishonesty in LLMs.
Acknowledgements
We would like to acknowledge Andy Zou for providing significant guidance for running the representation engineering experiments as well as Steven Basart for compute support.
@@ Appendix A Additional Results
A.1 Belief Consistency
@@
As discussed in Section 3, the standard definition of lying involves making a statement that one knows or believes to be false. If a model does not, in some sense, “believe” anything, it cannot lie under that definition. Therefore, measuring the model’s beliefs provides a basis for testing when it deliberately contradicts those beliefs.
@@
As described in Section 4.2, we elicit the model’s belief about a proposition using a neutral prompt. We repeat this question three times to verify consistency, then ask two additional prompts that probe related details of the proposition. If all these answers remain consistent, we label the belief as robustly held. In cases where models do not express consistent beliefs, we do not count the model as having lied. In Figure 11, we show that as LLMs increase in scale, their answers are marked as consistent beliefs by our pipeline more often. This aligns with prior findings that larger models maintain more coherent internal representations of factual information.
A.2 Honesty Utilities
@@
In Figure 12, we refer to $\Delta U$ as $U(\text{Honesty})-U(\text{Dishonesty})$ for clarity.
@@ Results.
Figure 12 shows the probability of lying on MASK examples plotted against $\Delta U$. We observe a negative correlation, meaning that models which value honesty more lie less often. In this experiment, we exclude examples where models do not have consistent beliefs when computing the percentage of cases where models lie, as in Appendix D. Including examples where models do not have consistent beliefs weakens the correlation to $-51.9\%$, indicating that the effect is still present across different variations of our metrics. This finding supports the idea that lying can be interpreted as utility maximization for models whose internal utilities do not sufficiently value honesty.
@@
Appendix B Data Collection
@@
B.1 Human Collected Data
@@
B.2 Automated Data Generation
@@
B.3 Final Manual Review
@@
Appendix C Consistently Honest Metric (Lying@$n$)
@@
We also present compute correlations for this metric in Figure 14, finding that the correlation is weaker, meaning that when they are honest, larger models are more consistently so.
@@
Appendix D Normalizing for Beliefs
We also consider a modified evaluation method which excludes examples for which the model being tested does not have a belief, as determined by our belief elicitation process. Given the results in Section A.1, one concern may be that the negative correlation between the honesty score and compute is simply due to smaller models not having beliefs about a given topic, as the honesty score includes honest responses, evasive responses, and cases where the model does not indicate a belief. Here, we “level the playing field” by removing the latter cases. Conceptually, the negative correlation in Figure 15 implies that as the models become larger and gain more knowledge, they opt to respond to the pressure prompts corresponding to this extra knowledge dishonestly more often than they opt to refuse or respond honestly.

### v3

@@ 4.3 Metrics
Accuracy measures a model’s truthfulness by evaluating the alignment between its belief $B$ and the ground truth $T$ for each proposition. We compute $\text{Accuracy}(B,T)=0$ if $B\neq T$, marking such cases as inaccurate. When a model shows no clear belief, we note this separately and exclude it from the final accuracy calculation. The final accuracy score is the average accuracy across all examples.
@@ 5 Experiments
We use MASK to explore two main experiments around model honesty. First, we conduct a comprehensive assessment of the degree of lying exhibited by state-of-the-art language models. Second, we evaluate the two baseline interventions designed to improve model honesty.
@@ 5.1 Models Evaluation
Figure˜6 reports the honesty rates of state-of-the-art models, which are quite low. No model shown is explicitly honest in more than $46\%$ of cases. GPT-4o and Llama-405B tell more lies than Claude 3.7 Sonnet, and most models are dishonest more than a third of the time. These lies appear even in short, straightforward scenarios, implying that current instruction tuning techniques alone are insufficient to prevent dishonesty. We also measure each model’s factual accuracy (Figure˜6) and observe that highly capable models tend to have over $70\%$ accuracy in their beliefs but do not necessarily exhibit higher honesty. Figure˜8 further shows that this finding holds across archetypes, indicating that these models can be pressured to lie in many different settings.
@@
In Figure˜7, we show that increased compute (FLOP) does not lead to more honest models, showing a negative correlation (Spearman coefficient: $59.9\%$). This indicates that higher scores are likely a result of design decisions and fine-tuning, rather than due solely to model capabilities obtained during pre-training. In contrast, accuracy strongly correlates with the amount of training compute used (Spearman coefficient: $87.3\%$). While model scaling improves factual accuracy, scaling AI models does not solve their tendency to produce intentional falsehoods when pressured.
For additional results, including a preliminary analysis of why different models have different honesty scores and a corroboration of our evaluation with self-reported lying, see Appendix˜A.
@@ 5.2 Improving the Honesty of Models
We test two baseline interventions to improve the honesty of two small Llama models and evaluate their effectiveness using MASK.
@@
For each pressure prompt, we prepend text to the system prompt, as if there were a separate “developer system prompt”, described in Section˜A.6. Although this approach improved the honesty of the responses in many examples ($12.2\%$ for Llama-2-7B, $8.8\%$ for Llama-2-13B), it still leaves substantial room for improvement (Figure˜9). This highlights the challenge that, in safety-critical domains, a model’s default behavior often requires more robust interventions than prompt engineering alone. Relying on specialized prompt engineering can also be tiresome; models should default to honest behavior without extensive developer prompt engineering.
@@
The technique is described further by Zou et al. (2023). While LoRRA led to measurable improvements in model honesty ($6.6\%$ for Llama-2-7B, $13.1\%$ for Llama-2-13B), it was also not sufficient to eliminate all dishonesty. This suggests that representation engineering may require methodological improvements to be fully robust in controlling lying in large language models.
@@ 6 Conclusion
Our dataset, MASK, along with our experiments, revealed that highly accurate LLMs will still engage in lies of commission. Scaling alone does not guarantee model honesty. Early fixes (targeted prompts and representation engineering) can help but remain imperfect, underscoring the need to define and study lying as a separate safety goal.
@@ Appendix A Additional Results
A.1 Broader Impacts
This work introduces the MASK benchmark and evaluation pipeline for measuring when large language models (LLMs) knowingly produce false statements under pressure. By publicly releasing 1,0001,\!000 carefully curated examples (with a private $500$-example hold-out set) that span misinformation, fabricated statistics, and other real-world deception scenarios, we aim to give developers a rigorous tool for diagnosing and reducing dishonest behavior and enable researchers to track progress on this safety-critical dimension. Positive impacts include fostering more transparent model comparisons and motivating techniques that measurably improve honesty. We do not believe there are any significant negative impacts of this work.
A.2 Limitations
While MASK offers a first-of-its-kind, large-scale benchmark for disentangling honesty from accuracy, it is far from comprehensive: (i) the 1,5001,\!500 examples focus on English textual prompts and short factual propositions, so the results may not generalize to multimodal inputs, longer dialogues, or non-English settings; (ii) our automated evaluation relies on an LLM judge whose mapping of responses to proposition values achieves $86.4\%$ agreement with human annotators, leaving non-trivial room for misclassification noise that could bias model rankings; (iii) the pressure-prompt archetypes target six hand-crafted scenarios and may miss other real-world incentives to deceive, such as multi-step planning or collusion between agents; (iv) because MASK tests models in isolation, it does not address interactive mitigations (e.g., tool use, chain-of-thought transparency, or external verification) that practitioners could employ, so reported honesty rates should be interpreted as worst-case propensities rather than deployment-time performance.
A.3 Belief Consistency
@@
As discussed in Section˜3, the standard definition of lying involves making a statement that one knows or believes to be false. If a model does not, in some sense, “believe” anything, it cannot lie under that definition. Therefore, measuring the model’s beliefs provides a basis for testing when it deliberately contradicts those beliefs.
@@
As described in Section˜4.2, we elicit the model’s belief about a proposition using a neutral prompt. We repeat this question three times to verify consistency, then ask two additional prompts that probe related details of the proposition. If all these answers remain consistent, we label the belief as robustly held. In cases where models do not express consistent beliefs, we do not count the model as having lied. In Figure˜10, we show that as LLMs increase in scale, their answers are marked as consistent beliefs by our pipeline more often. This aligns with prior findings that larger models maintain more coherent internal representations of factual information.
A.4 Honesty Utilities
@@
In Figure˜11, we refer to $\Delta U$ as $U(\text{Honesty})-U(\text{Dishonesty})$ for clarity.
@@ Results.
Figure˜11 shows the probability of lying on MASK examples plotted against $\Delta U$. We observe a negative correlation, meaning that models which value honesty more lie less often. In this experiment, we exclude examples where models do not have consistent beliefs when computing the percentage of cases where models lie, as in Section˜A.9. Including examples where models do not have consistent beliefs weakens the correlation to $51.9\%$, indicating that the effect is still present across different variations of our metrics. This finding supports the idea that lying can be interpreted as utility maximization for models whose internal utilities do not sufficiently value honesty.
@@
A.5 Self-Reported Lying
Experimental setup.
Results.
A.6 Developer System Prompt
USER SYSTEM PROMPT: <original_system_prompt>
A.7 Data Collection
@@
A.7.1 Human Collected Data
@@
A.7.2 Automated Data Generation
@@
A.7.3 Final Manual Review
@@
A.8 Consistently Honest Metric (Lying@$n$)
@@
We also present compute correlations for this metric in Figure˜14, finding that the correlation is weaker, meaning that when they are honest, larger models are more consistently so.
@@
A.9 Normalizing for Beliefs
We also consider a modified evaluation method which excludes examples for which the model being tested does not have a belief, as determined by our belief elicitation process. Given the results in Section˜A.3, one concern may be that the negative correlation between the honesty score and compute is simply due to smaller models not having beliefs about a given topic, as the honesty score includes honest responses, evasive responses, and cases where the model does not indicate a belief. Here, we “level the playing field” by removing the latter cases. Conceptually, the negative correlation in Figure˜15 implies that as the models become larger and gain more knowledge, they opt to respond to the pressure prompts corresponding to this extra knowledge dishonestly more often than they opt to refuse or respond honestly.